OpenStudy (anonymous):

A kids’ splash pool and an adults’ pool are shaped like similar triangles. The dimensions of the kids’ pool are one-fourth of the dimensions of the adults’ pool. If the area of the adults’ pool is 192 square meters, what is the area of the kids’ pool? Answer 10 m2 12 m2 24 m2 48 m2

OpenStudy (kira_yamato):

Given that they are similar ∆s \[\frac{A _{1}}{A _{2}} = \left[ \frac{l _{1}}{l _{2}} \right]^2\] A(kid pool) = (1/4)^2 * 192 = 1/16 * 192 = 12 sq. m